# Learnability in Online Kernel Selection with Memory Constraint via Data-dependent Regret Analysis

### v1

@@ 6 Conclusion
Learnability is an essential problem in online kernel selection with memory constraint. Previous work gave a negative result on learnability via worst-case regret analysis. We give positive results via data-dependent regret analysis. We characterize the regret bounds via two data-dependent complexities, i.e., kernel alignment and the cumulative losses of competitive hypothesis, and explicitly balance the data-dependent regret bounds and memory constraint. If kernel functions match well with the data, that is, the kernel alignment and the cumulative losses are sub-linear, then sub-linear regret can be achieved within a $\Theta(\ln{T})$ memory constraint and the corresponding hypothesis space is learnable.
Compared with the worst-case regret analysis which uses data size as data complexity, data-dependent regret analysis uses tighter data complexities to characterize the hardness of problems, and gives more insightful relationship among online learnability, memory constraint and data complexity. For the future work, it is interesting to study new data complexities under different loss functions.

### v2

@@ 7 Conclusion
Learnability is an essential problem in online kernel selection with memory constraint. Our work gave positive results on learnability through data-dependent regret analysis, in contrast to previous negative results obtained from worst-case regret analysis. We characterized the regret bounds via the kernel alignment and the cumulative losses, and gave new trade-offs between regret and memory constraint. If there is a kernel function matches well with the data, i.e., the kernel alignment and the cumulative losses is sub-linear, then sub-linear regret can be achieved within a $\Theta(\ln{T})$ memory constraint and the corresponding hypothesis space is learnable. Data-dependent regret analysis provides a new perspective for studying the learnability in online kernel selection with memory constraint.